OpenStudy (anonymous):

Solve the IVP ty′ + (t + 1)y = t, y(ln 2) = 1 for t > 0

myininaya (myininaya):

Divide both sides by t first: \[y'+\frac{t+1}{t}y=1\] now we should choose a v to multiply on both sides so that we can express the left hand side as vy'+v'y so we can write (vy)' so we need \[v'=\frac{t+1}{t}v\] but we can find this v by using separation of variables \[\frac{dv}{dt}=\frac{t+1}{t}v\] \[\frac{1}{v}dv=(1+\frac{1}{t}) dt\] now integrate both sides giving us \[lnv=t+lnt+C\] we don't need the constant (let C=0) \[lnv=t+lnt\] \[v=e^{t+lnt}\] so we multiply both sides by our v giving us: \[e^{t+lnt}y'+e^{t+lnt}*\frac{t+1}{t}y=e^{t+lnt}\] so now we can write: \[(e^{t+lnt}y)'=e^{t+lnt}\] now we integrate both sides because we know this will bring us closer to solving it: \[e^{t+lnt}y=\int\limits_{}^{}e^{t+lnt}dt+K\] but \[\int\limits_{}^{}e^{t+lnt}dt=\int\limits_{}^{}e^te^{lnt}dt=\int\limits_{}^{}te^tdt=te^t-\int\limits_{}^{}e^tdt=te^t-e^t+K_1\] so we have \[e^{t+lnt}y=te^t-e^t+K\] solving for y gives: \[y=\frac{te^t-e^t+K}{e^{t+lnt}}\]

myininaya (myininaya):

and that was for v>0 of course also you have initial condition y(ln2)=1 so just plug in ln2 and set =1 to solve for the constant K \[1=y(\ln2)=\frac{\ln2*e^{\ln2}-e^{\ln2}+K}{e^{\ln2+\ln(\ln2)}}\]

OpenStudy (anonymous):

I got y=1-1/t+2/te^t treating it like a linear equation.

myininaya (myininaya):

yes thats right we can simplify what i have gj dave

OpenStudy (anonymous):

Myininaya, can you help me with a de problem?

myininaya (myininaya):

\[y=\frac{te^t-e^t+k}{e^te^{lnt}}=\frac{t-1}{e^{lnt}}+\frac{K}{te^t}=\frac{t-1}{t}+\frac{K}{te^t}\] \[=1-\frac{1}{t}+\frac{K}{te^t}\]

myininaya (myininaya):

wendy are you there? i went the long way to solve it but there is a short way using a formula

OpenStudy (anonymous):

yes! I understand what you did to solve it. Thank you so much!!

myininaya (myininaya):

np